If a particle moves fron x to y in two different time interval of ratio 2 is to 1 with speed v1 and v2 then average dpeed if particle is

Grade:12th pass

1 Answers

John Sharon Sandeep
6 years ago
Hi Dear Buddy,Heres your answer..Let the distance between x and y be `S`Let T1= tTherefore T2=2t Applying speed=distance/time v1=S/T1 , v2=S/T2 Average velocity=(v1+v2)/2 =(S/T1 + S/T2) / 2 =(S/t + S/2t) / 2 =3S/4t.☺